# The N3LO Twist-2 Matching of Helicity TMDs and SIDIS $q_{\ast}$ Spectrum

### v3

@@ 4.4.2 Time-like results in $\overline{\text{MS}}$}
$\Delta P_{gg}^{T,\text{s}(2)}=C_{A}^{3}\bigg[\left(\left(-\frac{176}{3}\right)H_{-1,-1,0}+\frac{176}{3}H_{-1,0,1}\right)z^{2}+\left(336H_{-1,-1,0}+368H_{-1,0,1}\right)z+\frac{1}{z-1}\left(\left(-\frac{544}{3}\right)H_{1,0,0}+\left(-\frac{1072}{9}\right)\zeta_{2}+\left(-\frac{8}{9}\right)H_{1}+\frac{88}{3}\zeta_{3}+\frac{2920}{9}H_{1,0}+\frac{416\zeta_{2}}{3}H_{1}+192H_{1}\zeta_{3}+88\zeta_{4}-256\zeta_{2}H_{0,-1}-64\zeta_{2}H_{1,0}-208H_{0,0,1}-144H_{0,1,0}-256H_{0,0,1,1}-608H_{0,1,0,0}-256H_{0,1,0,1}-256H_{0,1,1,0}+128H_{1,0,-1,0}-640H_{1,0,0,0}-256H_{1,0,0,1}-256H_{1,0,1,0}-256H_{1,1,0,0}+\frac{490}{3}\right)+\frac{1}{z}\left(\left(-\frac{176}{3}\right)H_{-1,-1,0}+\frac{176}{3}H_{-1,0,1}\right)+\frac{z}{z-1}\left(\left(-\frac{48478}{27}\right)H_{0}+\left(-\frac{2968}{3}\right)\zeta_{3}+\left(-\frac{320}{3}\right)H_{0,1}+\left(-\frac{284}{3}\right)H_{0,0}+\left(-\frac{592}{9}\right)\zeta_{2}+\frac{320}{3}H_{0,-1,0}+\frac{280\zeta_{2}}{3}H_{0}+472\zeta_{4}-96\zeta_{2}H_{0,0}+160\zeta_{2}H_{0,1}-112H_{0,0,0}+64H_{0,-1,-1,0}-224H_{0,-1,0,0}-128H_{0,-1,0,1}-64H_{0,0,-1,0}+128H_{0,0,0,0}-160H_{0,0,0,1}-64H_{0,0,1,0}+\frac{1}{z+1}+\frac{46102}{9}\right)+\frac{z}{z-1}\left(\left(-\frac{1232}{3}\right)H_{1,0}+\frac{16}{9}H_{1}-336H_{1}\zeta_{2}-288H_{1}\zeta_{3}+416\zeta_{2}H_{0,-1}+96\zeta_{2}H_{1,0}+584H_{0,0,1}+320H_{0,1,0}+568H_{1,0,0}+384H_{0,0,1,1}+672H_{0,1,0,0}+384H_{0,1,0,1}+384H_{0,1,1,0}-192H_{1,0,-1,0}+960H_{1,0,0,0}+384H_{1,0,0,1}+384H_{1,0,1,0}+384H_{1,1,0,0}\right)+\frac{z^{2}}{z-1}\left(\left(-\frac{3232}{3}\right)H_{0,-1,0}+\left(-\frac{9136}{9}\right)\zeta_{2}+\left(-\frac{592}{9}\right)H_{0,1}+\frac{16964}{9}H_{0,0}+\frac{6736}{3}H_{0,0,0}+\frac{8990}{3}H_{0}+\left(-\frac{2024\zeta_{2}}{3}\right)H_{0}-800H_{0}\zeta_{3}-920\zeta_{3}-576\zeta_{4}-608\zeta_{2}H_{0,0}-32\zeta_{2}H_{0,1}+320H_{0,-1,-1,0}-544H_{0,-1,0,0}-704H_{0,0,-1,0}+1984H_{0,0,0,0}+800H_{0,0,0,1}+512H_{0,0,1,0}+\frac{1}{z+1}+\frac{46102}{9}\right)$ (144)
$\frac{z^{2}}{z-1}\left(\left(-\frac{544}{3}\right)H_{1,0,0}+\left(-\frac{8}{9}\right)H_{1}+\frac{176}{3}H_{0,1,0}+\frac{2920}{9}H_{1,0}+\frac{416\zeta_{2}}{3}H_{1}+192H_{1}\zeta_{3}-256\zeta_{2}H_{0,-1}-64\zeta_{2}H_{1,0}-112H_{0,0,1}-256H_{0,0,1,1}-288H_{0,1,0,0}-256H_{0,1,0,1}-256H_{0,1,1,0}+128H_{1,0,-1,0}-640H_{1,0,0,0}-256H_{1,0,0,1}-256H_{1,0,1,0}-256H_{1,1,0,0}\right)+\frac{z^{3}}{z-1}\left(\left(-\frac{176}{3}\right)H_{0,0,1}+\frac{88\zeta_{2}}{3}H_{1}\right)+\frac{z^{3}}{z-1}\left(\frac{1060}{9}H_{0,0}+\frac{3104}{9}H_{0,1}+\frac{2792}{3}\zeta_{3}+\frac{13402}{9}H_{0}+\frac{160\zeta_{2}}{3}H_{0}+304\zeta_{2}-680\zeta_{4}+96\zeta_{2}H_{0,0}-256\zeta_{2}H_{0,1}-48H_{0,-1,0}+464H_{0,0,0}+128H_{0,-1,0,0}+256H_{0,-1,0,1}-64H_{0,0,-1,0}-256H_{0,0,0,0}-96H_{0,0,0,1}-128H_{0,0,1,0}+\frac{1}{z+1}-\frac{47572}{9}\right)+\frac{z^{4}}{z-1}\left(\left(-\frac{176}{3}\right)H_{0,0,0}+\frac{176}{3}H_{0,-1,0}+\frac{440}{3}\zeta_{3}+\frac{352\zeta_{2}}{3}H_{0}+\frac{1}{z+1}\right)+\frac{1}{z+1}\left(\left(-\frac{6032}{9}\right)H_{-1,0}-288H_{-1}\zeta_{2}-384H_{-1}\zeta_{3}+512\zeta_{2}H_{-1,-1}-64\zeta_{2}H_{-1,0}-480H_{-1,0,0}+256H_{-1,-1,0,0}-512H_{-1,-1,0,1}+256H_{-1,0,-1,0}-640H_{-1,0,0,0}-128H_{-1,0,1,0}\right)+\frac{1}{(z-1)(z+1)}\left(\left(-\frac{89182}{27}\right)H_{0}+\left(-\frac{14612}{9}\right)H_{0,0}+\left(-\frac{4448}{3}\right)H_{0,0,0}+\frac{2144}{3}\zeta_{3}+\frac{2528}{3}H_{0,-1,0}+\frac{9136}{9}\zeta_{2}+528H_{0}\zeta_{2}+1088H_{0}\zeta_{3}+808\zeta_{4}+864\zeta_{2}H_{0,0}+304H_{0,1}-512H_{0,-1,-1,0}+768H_{0,-1,0,0}+960H_{0,0,-1,0}-2496H_{0,0,0,0}-1184H_{0,0,0,1}-768H_{0,0,1,0}-\frac{47572}{9}\right)+\frac{1}{z(z+1)}\left(\frac{176}{3}H_{-1,0,0}-88H_{-1}\zeta_{2}\right)+\frac{z}{z+1}\left(\left(-\frac{4736}{3}\right)H_{-1,0}-400H_{-1}\zeta_{2}-576H_{-1}\zeta_{3}+768\zeta_{2}H_{-1,-1}-96\zeta_{2}H_{-1,0}-1312H_{-1,0,0}+384H_{-1,-1,0,0}-768H_{-1,-1,0,1}+384H_{-1,0,-1,0}-960H_{-1,0,0,0}-192H_{-1,0,1,0}\right)+\frac{z^{2}}{z+1}\left(\left(-\frac{6032}{9}\right)H_{-1,0}-288H_{-1}\zeta_{2}-384H_{-1}\zeta_{3}+512\zeta_{2}H_{-1,-1}-64\zeta_{2}H_{-1,0}-480H_{-1,0,0}+256H_{-1,-1,0,0}-512H_{-1,-1,0,1}+256H_{-1,0,-1,0}-640H_{-1,0,0,0}-128H_{-1,0,1,0}\right)+\frac{z^{3}}{z+1}\left(\frac{176}{3}H_{-1,0,0}-88H_{-1}\zeta_{2}\right)+\frac{88\zeta_{2}}{3(z-1)z}H_{1}+336H_{-1,-1,0}+368H_{-1,0,1}\bigg]+N_{f}C_{A}^{2}\bigg[\left(\left(-\frac{32}{3}\right)H_{-1,-1,0}+\frac{32}{3}H_{-1,0,1}+\frac{16\zeta_{2}}{3}H_{1}-16H_{-1}\zeta_{2}\right)z^{2}+32H_{-1,0,1}+96H_{0,-1,-1,0}+\left(\left(-\frac{2174}{9}\right)H_{1}+56H_{-1}\zeta_{2}+88H_{1}\zeta_{2}-48\zeta_{2}H_{0,-1}+176H_{-1,-1,0}+32H_{-1,0,1}-96H_{0,-1,-1,0}+48H_{0,-1,0,0}+80H_{0,0,-1,0}+24H_{0,0,0,1}+24H_{0,1,0,0}\right)z+\frac{2174}{9}H_{1}-48H_{0,-1,0,0}-16H_{0,0,-1,0}+\frac{1}{z-1}\left(\left(-\frac{1844}{9}\right)H_{0,1}+\left(-\frac{356}{3}\right)H_{1,0,0}+\left(-\frac{928}{9}\right)H_{1,0}+\left(-\frac{112}{3}\right)\zeta_{3}+\frac{160}{9}\zeta_{2}+48\zeta_{2}H_{0,1}-124H_{0,0,1}-96H_{0,1,0}-\frac{836}{27}\right)+16H_{0,0,0,0}+24H_{0,0,0,1}+24H_{0,1,0,0}+\frac{1}{z}\left(\left(-\frac{32}{3}\right)H_{-1,-1,0}+\frac{32}{3}H_{-1,0,1}+\left(-\frac{16\zeta_{2}}{3}\right)H_{1}-16H_{-1}\zeta_{2}\right)+\frac{z}{z-1}\left(\left(-\frac{1396}{9}\right)H_{0,0}+\left(-\frac{368}{3}\right)\zeta_{3}+\left(-\frac{1138}{27}\right)H_{0}+\frac{376}{3}H_{0,-1,0}+\frac{508}{3}\zeta_{2}+\left(-\frac{100\zeta_{2}}{3}\right)H_{0}-138\zeta_{4}+24\zeta_{2}H_{0,0}-32H_{0,0,0}+\frac{1}{z+1}+\frac{23740}{27}\right)+56H_{-1}\zeta_{2}-88H_{1}\zeta_{2}+48\zeta_{2}H_{0,-1}+176H_{-1,-1,0}$ (145)
[…]

### v4

@@ 4.4.2 Time-like results in $\overline{\text{MS}}$}
$\Delta P_{gg}^{T,\text{s}(2)}=C_{A}^{3}\bigg[\left(\left(-\frac{176}{3}\right)H_{-1,-1,0}+\frac{176}{3}H_{-1,0,1}\right)z^{2}+\left(336H_{-1,-1,0}+368H_{-1,0,1}\right)z+\frac{1}{z-1}\left(\left(-\frac{544}{3}\right)H_{1,0,0}+\left(-\frac{1072}{9}\right)\zeta_{2}+\left(-\frac{8}{9}\right)H_{1}+\frac{88}{3}\zeta_{3}+\frac{2920}{9}H_{1,0}+\frac{416\zeta_{2}}{3}H_{1}+192H_{1}\zeta_{3}+88\zeta_{4}-256\zeta_{2}H_{0,-1}-64\zeta_{2}H_{1,0}-208H_{0,0,1}-144H_{0,1,0}-256H_{0,0,1,1}-608H_{0,1,0,0}-256H_{0,1,0,1}-256H_{0,1,1,0}+128H_{1,0,-1,0}-640H_{1,0,0,0}-256H_{1,0,0,1}-256H_{1,0,1,0}-256H_{1,1,0,0}+\frac{490}{3}\right)+\frac{1}{z}\left(\left(-\frac{176}{3}\right)H_{-1,-1,0}+\frac{176}{3}H_{-1,0,1}\right)+\frac{z}{z-1}\left(\left(-\frac{48478}{27}\right)H_{0}+\left(-\frac{2968}{3}\right)\zeta_{3}+\left(-\frac{320}{3}\right)H_{0,1}+\left(-\frac{284}{3}\right)H_{0,0}+\left(-\frac{592}{9}\right)\zeta_{2}+\frac{320}{3}H_{0,-1,0}+\frac{280\zeta_{2}}{3}H_{0}+472\zeta_{4}-96\zeta_{2}H_{0,0}+160\zeta_{2}H_{0,1}-112H_{0,0,0}+64H_{0,-1,-1,0}-224H_{0,-1,0,0}-128H_{0,-1,0,1}-64H_{0,0,-1,0}+128H_{0,0,0,0}-160H_{0,0,0,1}-64H_{0,0,1,0}+\frac{1}{z+1}+\frac{46102}{9}\right)+\frac{z}{z-1}\left(\left(-\frac{1232}{3}\right)H_{1,0}+\frac{16}{9}H_{1}-336H_{1}\zeta_{2}-288H_{1}\zeta_{3}+416\zeta_{2}H_{0,-1}+96\zeta_{2}H_{1,0}+584H_{0,0,1}+320H_{0,1,0}+568H_{1,0,0}+384H_{0,0,1,1}+672H_{0,1,0,0}+384H_{0,1,0,1}+384H_{0,1,1,0}-192H_{1,0,-1,0}+960H_{1,0,0,0}+384H_{1,0,0,1}+384H_{1,0,1,0}+384H_{1,1,0,0}\right)+\frac{z^{2}}{z-1}\left(\left(-\frac{3232}{3}\right)H_{0,-1,0}+\left(-\frac{9136}{9}\right)\zeta_{2}+\left(-\frac{592}{9}\right)H_{0,1}+\frac{16964}{9}H_{0,0}+\frac{6736}{3}H_{0,0,0}+\frac{8990}{3}H_{0}+\left(-\frac{2024\zeta_{2}}{3}\right)H_{0}-800H_{0}\zeta_{3}-920\zeta_{3}-576\zeta_{4}-608\zeta_{2}H_{0,0}-32\zeta_{2}H_{0,1}+320H_{0,-1,-1,0}-544H_{0,-1,0,0}-704H_{0,0,-1,0}+1984H_{0,0,0,0}+800H_{0,0,0,1}+512H_{0,0,1,0}+\frac{1}{z+1}+\frac{46102}{9}\right)$
$\left.+\frac{z^{4}}{z-1}\bigg(-\frac{176}{3}H_{0,0,0}+\frac{176}{3}H_{0,-1,0}+\frac{440}{3}\zeta_{3}+\frac{352\zeta_{2}}{3}H_{0}+\frac{1}{z+1}\bigg)\right.+\frac{z^{2}}{z-1}\bigg(-\frac{544}{3}H_{1,0,0}-\frac{8}{9}H_{1}+\frac{176}{3}H_{0,1,0}+\frac{2920}{9}H_{1,0}+\frac{416\zeta_{2}}{3}H_{1}+192H_{1}\zeta_{3}-256\zeta_{2}H_{0,-1}-64\zeta_{2}H_{1,0}-112H_{0,0,1}-256H_{0,0,1,1}-288H_{0,1,0,0}-256H_{0,1,0,1}-256H_{0,1,1,0}+128H_{1,0,-1,0}-640H_{1,0,0,0}-256H_{1,0,0,1}-256H_{1,0,1,0}-256H_{1,1,0,0}\bigg)+\frac{z^{3}}{z-1}\bigg(-\frac{176}{3}H_{0,0,1}+\frac{88\zeta_{2}}{3}H_{1}\bigg)+\frac{z^{3}}{z-1}\bigg(\frac{1060}{9}H_{0,0}+\frac{3104}{9}H_{0,1}+\frac{2792}{3}\zeta_{3}+\frac{13402}{9}H_{0}+\frac{160\zeta_{2}}{3}H_{0}+304\zeta_{2}-680\zeta_{4}+96\zeta_{2}H_{0,0}-256\zeta_{2}H_{0,1}-48H_{0,-1,0}+464H_{0,0,0}+128H_{0,-1,0,0}+256H_{0,-1,0,1}-64H_{0,0,-1,0}-256H_{0,0,0,0}-96H_{0,0,0,1}-128H_{0,0,1,0}+\frac{1}{z+1}-\frac{47572}{9}\bigg)+\frac{1}{z+1}\bigg(-\frac{6032}{9}H_{-1,0}-288H_{-1}\zeta_{2}-384H_{-1}\zeta_{3}+512\zeta_{2}H_{-1,-1}-64\zeta_{2}H_{-1,0}-480H_{-1,0,0}+256H_{-1,-1,0,0}-512H_{-1,-1,0,1}+256H_{-1,0,-1,0}-640H_{-1,0,0,0}-128H_{-1,0,1,0}\bigg)+\frac{1}{(z-1)(z+1)}\bigg(-\frac{89182}{27}H_{0}-\frac{14612}{9}H_{0,0}-\frac{4448}{3}H_{0,0,0}+\frac{2144}{3}\zeta_{3}+\frac{2528}{3}H_{0,-1,0}+\frac{9136}{9}\zeta_{2}+528H_{0}\zeta_{2}+1088H_{0}\zeta_{3}+808\zeta_{4}+864\zeta_{2}H_{0,0}+304H_{0,1}-512H_{0,-1,-1,0}+768H_{0,-1,0,0}+960H_{0,0,-1,0}-2496H_{0,0,0,0}-1184H_{0,0,0,1}-768H_{0,0,1,0}-\frac{47572}{9}\bigg)+\frac{1}{z(z+1)}\left(\frac{176}{3}H_{-1,0,0}-88H_{-1}\zeta_{2}\right)+\frac{z}{z+1}\bigg(-\frac{4736}{3}H_{-1,0}-400H_{-1}\zeta_{2}-576H_{-1}\zeta_{3}+768\zeta_{2}H_{-1,-1}-96\zeta_{2}H_{-1,0}-1312H_{-1,0,0}+384H_{-1,-1,0,0}-768H_{-1,-1,0,1}+384H_{-1,0,-1,0}-960H_{-1,0,0,0}-192H_{-1,0,1,0}\bigg)+\frac{z^{2}}{z+1}\bigg(-\frac{6032}{9}H_{-1,0}-288H_{-1}\zeta_{2}-384H_{-1}\zeta_{3}+512\zeta_{2}H_{-1,-1}-64\zeta_{2}H_{-1,0}-480H_{-1,0,0}+256H_{-1,-1,0,0}-512H_{-1,-1,0,1}+256H_{-1,0,-1,0}-640H_{-1,0,0,0}-128H_{-1,0,1,0}\bigg)+\frac{z^{3}}{z+1}\left(\frac{176}{3}H_{-1,0,0}-88H_{-1}\zeta_{2}\right)+\frac{88\zeta_{2}}{3(z-1)z}H_{1}+336H_{-1,-1,0}+368H_{-1,0,1}\bigg]+N_{f}C_{A}^{2}\bigg[\left(-\frac{32}{3}H_{-1,-1,0}+\frac{32}{3}H_{-1,0,1}+\frac{16\zeta_{2}}{3}H_{1}-16H_{-1}\zeta_{2}\right)z^{2}+32H_{-1,0,1}+96H_{0,-1,-1,0}+\left(-\frac{2174}{9}H_{1}+56H_{-1}\zeta_{2}+88H_{1}\zeta_{2}-48\zeta_{2}H_{0,-1}+176H_{-1,-1,0}+32H_{-1,0,1}-96H_{0,-1,-1,0}+48H_{0,-1,0,0}+80H_{0,0,-1,0}+24H_{0,0,0,1}+24H_{0,1,0,0}\right)z+\frac{2174}{9}H_{1}-48H_{0,-1,0,0}-16H_{0,0,-1,0}+\frac{1}{z-1}\bigg(-\frac{1844}{9}H_{0,1}-\frac{356}{3}H_{1,0,0}-\frac{928}{9}H_{1,0}-\frac{112}{3}\zeta_{3}+\frac{160}{9}\zeta_{2}+48\zeta_{2}H_{0,1}-124H_{0,0,1}-96H_{0,1,0}-\frac{836}{27}\bigg)+16H_{0,0,0,0}+24H_{0,0,0,1}+24H_{0,1,0,0}+\frac{1}{z}\bigg(-\frac{32}{3}H_{-1,-1,0}+\frac{32}{3}H_{-1,0,1}+\left(-\frac{16\zeta_{2}}{3}\right)H_{1}-16H_{-1}\zeta_{2}\bigg)+\frac{z}{z-1}\bigg(-\frac{1396}{9}H_{0,0}-\frac{368}{3}\zeta_{3}-\frac{1138}{27}H_{0}+\frac{376}{3}H_{0,-1,0}+\frac{508}{3}\zeta_{2}+\left(-\frac{100\zeta_{2}}{3}\right)H_{0}-138\zeta_{4}+24\zeta_{2}H_{0,0}-32H_{0,0,0}+\frac{1}{z+1}+\frac{23740}{27}\bigg)+56H_{-1}\zeta_{2}-88H_{1}\zeta_{2}+48\zeta_{2}H_{0,-1}+176H_{-1,-1,0}+\frac{z}{z-1}\left(\frac{512}{3}H_{0,1}+\frac{512}{3}H_{1,0}+200H_{0,0,1}+128H_{0,1,0}+200H_{1,0,0}\right)$
[…]